# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 4.2. Open Kulikov surfaces from polarized log Calabi-Yau surfaces
Using the deformation theory results proved in §4.1, we will glue these local $\mathbb{Q}$-Gorenstein deformations of the germs $(V_{ij},\partial V_{ij})$ into a global deformation of $(\widetilde{X}_{\mathscr{P}}^{v},\partial\widetilde{X}_{\mathscr{P}}^{v})$. Let $Q_{v}$ be the number triangles $\Delta_{ij}$ with $pr(p_{ij})=v$. As a projective toric variety $X$ always satisfies $H^{1}(X,\mathcal{O}_{X})=0$, it follows from Lemma 4.2 that there exists a $Q_{v}$-dimensional $\mathbb{Q}$-Gorenstein deformation of $(\widetilde{X}_{\mathscr{P}}^{v},\partial\widetilde{X}_{\mathscr{P}}^{v})$ induced by the one-dimensional $\mathbb{Q}$-Gorenstein deformations of the $Q_{v}$ germs $(V_{ij},\partial V_{ij})$. Moreover, the base of this $Q_{v}$-dimensional deformation is irreducible, since it is the case for each of the one-dimensional deformations of the germs $(V_{ij},\partial V_{ij})$. We denote by $(X_{\mathscr{P}}^{v},\partial X_{\mathscr{P}}^{v})$ a general fiber of this deformation. It is a log Calabi–Yau surface of charge $Q_{v}$, well-defined up to smooth deformations since the base of the deformation is irreducible.
STEP III: In a third step we glue together the log Calabi-Yau surfaces $X_{\mathscr{P}}^{v}$ for $v\in\mathscr{P}^{[0]}$ obtained in Step II along the divisors corresponding to the edges of $\mathscr{P}$, to obtain the normal crossing surface $\mathcal{X}_{\mathscr{P},0}$. Note that such a gluing is not unique: if such an edge lies in the interior of $P$, the corresponding divisor is a $\mathbb{P}^{1}$, and if it is on the boundary we have an $\mathbb{A}^{1}$. Therefore, the set of all possible gluings of the surfaces $X^{v}_{\mathscr{P}}$, for $v\in\mathscr{P}^{[0]}$, along a common divisor is a torsor for the torus (*$(\mathbb{C}^{*})^{|\mathscr{P}^{1}|}$, where $|\mathscr{P}^{1}|$ is the number of edges of the polyhedral decomposition $\mathscr{P}$ on $P$. Each $\mathbb{C}^{\star}$ acts by scaling of the gluing along the corresponding copy of $\mathbb{P}^{1}$ or $\mathbb{A}^{1}$.
@@
Now that we obtained $\mathcal{X}_{\mathscr{P},0}$, we investigate its properties below. In particular, we show that it is a quasi-projective open Kulikov surface. To do that, we first describe the irreducible components $X^{v}_{\mathscr{P}}$ of $\mathcal{X}_{\mathscr{P},0}$ for $v\in(\partial P)_{\mathbb{Z}}$. The boundary $\partial P$ of $P$ is naturally identified with the intersection complex of the divisor $D$ polarized by the restriction $L|_{D}$ of $L$ to $D$. As $D$ is either a cycle of rational curves or a nodal curve of arithmetic genus one, it follows that $D$ is topologically a circle, and is of integral length $\mathrm{deg}L|_{D}$. Moreover, there are integral points $v_{x}\in(\partial P)_{\mathbb{Z}}$ in one-to-one correspondence with the 0-dimensional strata $x$ of $D$. If $x$ and $y$ are two 0-dimensional strata on the same irreducible component $D_{i}$ of $D$, then there is a corresponding interval $I_{D_{i}}\subset\partial P\simeq S^{1}$ with endpoints $v_{x}$ and $v_{y}$, of integral length $\mathrm{deg}L|_{D_{i}}$, and containing no other vertex of the form $v_{z}$.
For every $v\in(\partial P)_{\mathbb{Z}}$, we denote by
@@
the affinization of $(X^{v}_{\mathscr{P}},\partial X^{v}_{\mathscr{P}})$. In the following Lemmas 4.3-4.5, we give an explicit description of the affine schemes $\overline{X}^{v}_{\mathscr{P}}$ and of the corresponding affinization map $X^{v}_{\mathscr{P}}\rightarrow\overline{X}^{v}_{\mathscr{P}}$, which shows in particular that the algebras of regular functions $H^{0}(X^{v}_{\mathscr{P}},\mathcal{O}_{X^{v}_{\mathscr{P}}})$ and $H^{0}(\partial X^{v}_{\mathscr{P}},\mathcal{O}_{\partial X^{v}_{\mathscr{P}}})$ are finitely generated. To do that, note that by Proposition 2.1, for every 0-dimensional stratum $x$ of $D$, the germ of $(Y,D)$ at $x$ is a cyclic quotient surface singularity, which can be described torically by a fan consisting of a single rational polyhedral cone $\sigma_{x}\subset\mathbb{R}^{2}$.
@@
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface, with a polyhedral decomposition $\mathscr{P}$ of a Symington polytope $P$ defined as above. For every 00-dimensional stratum of $D$, $\overline{X}^{v_{x}}_{\mathscr{P}}$ is an affine surface obtained by a flat deformation of the affine toric surface $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$, and $\partial\overline{X}^{v_{x}}_{\mathscr{P}}$ is a flat locally trivial deformation of the union of the two toric divisors of $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$. Moreover, the affinization map $X^{v_{x}}_{\mathscr{P}}\rightarrow\overline{X}^{v_{x}}_{\mathscr{P}}$ is a surjective projective birational morphism.
@@
Assume first that there is no interior vertex $p_{ij}$ of a triangle $\Delta_{ij}$ as in §2.2 such that $pr(p_{ij})=v_{x}$. Then, by Step I of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, $X^{v_{x}}_{\mathscr{P}}$ is the toric surface whose fan is given by the star of $\mathscr{P}$ at $v_{x}$. Moreover, this fan is naturally a refinement of the strictly convex cone $\sigma_{x}$ giving a local description of $P$ near $v_{x}$. It follows that in this case the affinization $\overline{\mathcal{X}}_{\mathscr{P},0}$ is exactly the affine toric surface $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$, and the affinization map is a toric morphism induced by a refinement of fans, and so is surjective, projective, and birational.
@@
In general, if there are points $p_{ij}$ such that $pr(p_{ij})=v_{x}$, then, by Step II of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, $X^{v_{x}}_{\mathscr{P}}$ is a flat deformation, obtained by partial smoothing of the toric boundary, of the toric surface $X^{v_{x}}_{\overline{\mathscr{P}}}$ whose fan is given by the star of $\mathscr{P}$ at $v_{x}$, and whose affinization is given by $\overline{X}^{v_{x}}_{\overline{\mathscr{P}}}=\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$. As $X^{v_{x}}_{\overline{\mathscr{P}}}$ is a toric surface, we have $H^{1}(X^{v_{x}}_{\overline{\mathscr{P}}},\mathcal{O}_{X^{v_{x}}_{\overline{\mathscr{P}}}})=0$. Hence, by [96, Theorem 1.4] – see also [26, Theorem 3.1], the affinization $X^{v_{x}}_{\mathscr{P}}\rightarrow\overline{X}^{v_{x}}_{\mathscr{P}}$ is a flat deformation of the toric affinization $X^{v_{x}}_{\overline{\mathscr{P}}}\rightarrow\overline{X}^{v_{x}}_{\overline{\mathscr{P}}}=\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$, and this implies Lemma 4.3. ∎
@@
The affine toric surface $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$ is the dual quotient singularity of the quotient singularity $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}^{\vee}\cap\mathbb{Z}^{2}]$ describing the germ of $(Y,D)$ at $x$.
Lemma 4.5.
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface, with a polyhedral decomposition $\mathscr{P}$ of a Symington polytope $P$ defined as above. For every vertex $v\in(\partial P)_{\mathbb{Z}}$ such that $v\neq v_{x}$ for all 00-dimensional strata $x$ of $D$, we have $\overline{X}^{v}_{\mathscr{P}}\simeq\mathbb{A}^{1}$, and the affinization map $X^{v}_{\mathscr{P}}\rightarrow\overline{X}^{v}_{\mathscr{P}}\simeq\mathbb{A}^{1}$ is a surjective projective morphism with general fiber $\mathbb{P}^{1}$.
Proof.
Assume first that there is no interior vertex $p_{ij}$ of a triangle $\Delta_{ij}$ as in §2.2 such that $pr(p_{ij})=v$. Then, by Step I of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, $X^{v}_{\mathscr{P}}$ is the toric surface whose fan is given by the star of $\mathscr{P}$ at $v$. Since $v\neq v_{x}$ for all 00-dimensional strata $x$ of $D$, the boundary $\partial P$ of $P$ is straight at $v$, and so the fan of $X^{v}_{\mathscr{P}}$ contains two opposite rays coming from the two edges of $\partial P$ containing $v$, and all the other rays are contained in a half-plane delimited by these two opposite rays. It follows that there exists a surjective projective toric morphism $X^{v}_{\mathscr{P}}\rightarrow\mathbb{A}^{1}$ with general fiber $\mathbb{P}^{1}$. In particular, the affinization $\overline{X}^{v}_{\mathscr{P}}$ of $X^{v}_{\mathscr{P}}$ is isomorphic to $\mathbb{A}^{1}$, and $X^{v}_{\mathscr{P}}\rightarrow\mathbb{A}^{1}$ is the affinization morphism.
In general, if there are points $p_{ij}$ such that $pr(p_{ij})=v$, then, by Step II of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, $X^{v}_{\mathscr{P}}$ is a flat deformation, obtained by partial smoothing of the toric boundary, of the toric surface $X^{v}_{\overline{\mathscr{P}}}$ whose fan is given by the star of $\mathscr{P}$ at $v$. As in the previous case, we have a surjective projective toric morphism $X^{v}_{\overline{\mathscr{P}}}\rightarrow\mathbb{A}^{1}$ with general fiber $\mathbb{P}^{1}$. Since $X^{v}_{\overline{\mathscr{P}}}$ is a toric surface, we have $H^{1}(X^{v}_{\overline{\mathscr{P}}},\mathcal{O}_{X^{v}_{\overline{\mathscr{P}}}})=0$, and so by [96, Theorem 1.4], the morphism $X^{v}_{\overline{\mathscr{P}}}\rightarrow\mathbb{A}^{1}$ deforms to a surjective projective morphism $X^{v}_{\mathscr{P}}\rightarrow\mathbb{A}^{1}$ with general fiber $\mathbb{P}^{1}$. This morphism is necessarily the affinization map of $X^{v}_{\mathscr{P}}$, and so Lemma 4.3 follows. ∎

### v3

@@ 4.2. Open Kulikov surfaces from polarized log Calabi–Yau surfaces
Using the deformation theory results proved in §4.1, we will glue these local $\mathbb{Q}$-Gorenstein deformations of the germs $(V_{ij},\partial V_{ij})$ into a global deformation of $(\widetilde{X}_{\mathscr{P}}^{v},\partial\widetilde{X}_{\mathscr{P}}^{v})$. Let $Q_{v}$ be the number triangles $\Delta_{ij}$ with $pr(p_{ij})=v$. Since a projective toric variety $X$ always satisfies $H^{1}(X,\mathcal{O}_{X})=0$, we obtain by Lemma 4.2 that there exists a $Q_{v}$-dimensional $\mathbb{Q}$-Gorenstein deformation of $(\widetilde{X}_{\mathscr{P}}^{v},\partial\widetilde{X}_{\mathscr{P}}^{v})$ induced by the one-dimensional $\mathbb{Q}$-Gorenstein deformations of the $Q_{v}$ germs $(V_{ij},\partial V_{ij})$. Moreover, the base of this $Q_{v}$-dimensional deformation is irreducible, since it is the case for each of the one-dimensional deformations of the germs $(V_{ij},\partial V_{ij})$. We denote by $(X_{\mathscr{P}}^{v},\partial X_{\mathscr{P}}^{v})$ a general fiber of this deformation. It is a smooth, possibly non-compact, log Calabi–Yau surface of charge $Q_{v}$, well-defined up to smooth deformations since the base of the deformation is irreducible. If $v$ is an interior vertex of $P$, then $(X_{\mathscr{P}}^{v},\partial X_{\mathscr{P}}^{v})$ is a smooth projective log Calabi–Yau surface, that is, a $2$-surface as in Definition 3.6.
@@
If $v$ is a boundary vertex of $P$, then $(X_{\mathscr{P}}^{v},\partial X_{\mathscr{P}}^{v})$ is a smooth non-compact log Calabi–Yau surface, described explicitly in Lemmas 4.4-4.5 below. The boundary $\partial P$ of $P$ is naturally identified with the intersection complex of the divisor $D$ polarized by the restriction $L|_{D}$ of $L$ to $D$. As $D$ is either a cycle of rational curves or a nodal curve of arithmetic genus one, it follows that $\partial P$ is topologically a circle, and is of integral length $\mathrm{deg}L|_{D}$. Moreover, there are integral points $v_{x}\in(\partial P)_{\mathbb{Z}}$ in one-to-one correspondence with the 0-dimensional strata $x$ of $D$. If $x$ and $y$ are two 0-dimensional strata on the same irreducible component $D_{i}$ of $D$, then there is a corresponding interval $I_{D_{i}}\subset\partial P\simeq S^{1}$ with endpoints $v_{x}$ and $v_{y}$, of integral length $\mathrm{deg}L|_{D_{i}}$, and containing no other vertex of the form $v_{z}$. For every $v\in(\partial P)_{\mathbb{Z}}$, we denote by
@@
the affinization of $(X^{v}_{\mathscr{P}},\partial X^{v}_{\mathscr{P}})$. By Proposition 2.1, the germ of $(Y,D)$ at every 0-dimensional stratum $x$ of $D$ is a cyclic quotient surface singularity, which can be described torically by a fan consisting of a single rational polyhedral cone $\sigma_{x}\subset\mathbb{R}^{2}$. The affine toric surface $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$ is the dual quotient singularity of the quotient singularity $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}^{\vee}\cap\mathbb{Z}^{2}]$ describing the germ of $(Y,D)$ at $x$.
@@
Let $(Y,D,L)$ be a generic polarized log Calabi–Yau surface, with a polyhedral decomposition $\mathscr{P}$ of a Symington polytope $P$ defined as above. For every 0-dimensional stratum $x$ of $D$, $\overline{X}^{v_{x}}_{\mathscr{P}}$ is an affine surface obtained by a flat deformation of the affine toric surface $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$, and $\partial\overline{X}^{v_{x}}_{\mathscr{P}}$ is a flat locally trivial deformation of the union of the two toric divisors of $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$. Moreover, the affinization map $X^{v_{x}}_{\mathscr{P}}\rightarrow\overline{X}^{v_{x}}_{\mathscr{P}}$ is a surjective projective birational morphism, and every exceptional curve is contracted to the 0-dimensional stratum of $\partial\overline{X}^{v_{x}}_{\mathscr{P}}$. In particular, $(X^{v_{x}}_{\mathscr{P}},\partial X^{v_{x}}_{\mathscr{P}})$ is a 0-surface as in Definition 3.2.
@@
Assume first that there is no interior vertex $p_{ij}$ of a triangle $\Delta_{ij}$ as in §2.2 such that $pr(p_{ij})=v_{x}$. Then, by Step I, $X^{v_{x}}_{\mathscr{P}}$ is the toric surface whose fan is given by the star of $\mathscr{P}$ at $v_{x}$. Moreover, this fan is naturally a refinement of the strictly convex cone $\sigma_{x}$ giving a local description of $P$ near $v_{x}$. It follows that in this case the affinization $\mathcal{X}^{\mathrm{can}}_{\mathscr{P},0}$ is exactly the affine toric surface $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$, and the affinization map is a toric morphism induced by a refinement of fans, and so is surjective, projective, and birational.
@@
In general, if there are points $p_{ij}$ such that $pr(p_{ij})=v_{x}$, then, by Step II of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, $X^{v_{x}}_{\mathscr{P}}$ is a flat deformation, obtained by partial smoothing of the toric boundary, of the toric surface $X^{v_{x}}_{\overline{\mathscr{P}}}$ whose fan is given by the star of $\mathscr{P}$ at $v_{x}$, and whose affinization is given by $\overline{X}^{v_{x}}_{\overline{\mathscr{P}}}=\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$. Since $X^{v_{x}}_{\overline{\mathscr{P}}}$ is a toric surface, we have $H^{1}(X^{v_{x}}_{\overline{\mathscr{P}}},\mathcal{O}_{X^{v_{x}}_{\overline{\mathscr{P}}}})=0$. Hence, by [100, Theorem 1.4] – see also [27, Theorem 3.1], the affinization $X^{v_{x}}_{\mathscr{P}}\rightarrow\overline{X}^{v_{x}}_{\mathscr{P}}$ is a flat deformation of the toric affinization $X^{v_{x}}_{\overline{\mathscr{P}}}\rightarrow\overline{X}^{v_{x}}_{\overline{\mathscr{P}}}=\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$. Finally, every exceptional curve is contracted to the 0-dimensional stratum of $\partial\overline{X}^{v_{x}}_{\mathscr{P}}$ by Lemma 2.12. ∎
Lemma 4.5.
Let $(Y,D,L)$ be a generic polarized log Calabi–Yau surface, with a polyhedral decomposition $\mathscr{P}$ of a Symington polytope $P$ defined as above. For every vertex $v\in(\partial P)_{\mathbb{Z}}$ such that $v\neq v_{x}$ for all 0-dimensional strata $x$ of $D$, we have $\overline{X}^{v}_{\mathscr{P}}\simeq\mathbb{A}^{1}$. Moreover, the affinization map $X^{v}_{\mathscr{P}}\rightarrow\overline{X}^{v}_{\mathscr{P}}\simeq\mathbb{A}^{1}$ is a surjective projective morphism with general fiber $\mathbb{P}^{1}$, and every exceptional curve is contracted to $0\in\mathbb{A}^{1}$. In particular, $(X^{v_{x}}_{\mathscr{P}},\partial X^{v_{x}}_{\mathscr{P}})$ is a $1$-surface as in Definition 3.4.
Proof.
Assume first that there is no interior vertex $p_{ij}$ of a triangle $\Delta_{ij}$ as in §2.2 such that $pr(p_{ij})=v$. Then, by Step I of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, $X^{v}_{\mathscr{P}}$ is the toric surface whose fan is given by the star of $\mathscr{P}$ at $v$. Since $v\neq v_{x}$ for all 0-dimensional strata $x$ of $D$, the boundary $\partial P$ of $P$ is straight at $v$, and so the fan of $X^{v}_{\mathscr{P}}$ contains two opposite rays coming from the two edges of $\partial P$ containing $v$, and all the other rays are contained in a half-plane delimited by these two opposite rays. It follows that there exists a surjective projective toric morphism $X^{v}_{\mathscr{P}}\rightarrow\mathbb{A}^{1}$ with general fiber $\mathbb{P}^{1}$. In particular, the affinization $\overline{X}^{v}_{\mathscr{P}}$ of $X^{v}_{\mathscr{P}}$ is isomorphic to $\mathbb{A}^{1}$, and $X^{v}_{\mathscr{P}}\rightarrow\mathbb{A}^{1}$ is the affinization morphism.
In general, if there are points $p_{ij}$ such that $pr(p_{ij})=v$, then, $X^{v}_{\mathscr{P}}$ is a flat deformation, obtained by partial smoothing of the toric boundary, of the toric surface $X^{v}_{\overline{\mathscr{P}}}$ whose fan is given by the star of $\mathscr{P}$ at $v$. As in the previous case, we have a surjective projective toric morphism $X^{v}_{\overline{\mathscr{P}}}\rightarrow\mathbb{A}^{1}$ with general fiber $\mathbb{P}^{1}$. Since $X^{v}_{\overline{\mathscr{P}}}$ is a toric surface, we have $H^{1}(X^{v}_{\overline{\mathscr{P}}},\mathcal{O}_{X^{v}_{\overline{\mathscr{P}}}})=0$, and so by [100, Theorem 1.4], the morphism $X^{v}_{\overline{\mathscr{P}}}\rightarrow\mathbb{A}^{1}$ deforms to a surjective projective morphism $X^{v}_{\mathscr{P}}\rightarrow\mathbb{A}^{1}$ with general fiber $\mathbb{P}^{1}$. This morphism is necessarily the affinization map of $X^{v}_{\mathscr{P}}$. Finally, every exceptional curve is contracted to $0\in\mathbb{A}^{1}$ by Lemma 2.12. ∎
STEP III: In a third step we glue together the log Calabi–Yau surfaces $X_{\mathscr{P}}^{v}$ for $v\in\mathscr{P}^{[0]}$ obtained in Step II along the divisors corresponding to the edges of $\mathscr{P}$, to obtain a normal crossing surface $\mathcal{X}_{\mathscr{P},0}$, well-defined up to locally trivial deformations. We show that the normal crossing surface $\mathcal{X}_{\mathscr{P},0}$ is a quasi-projective open Kulikov surface as in Definition 3.7.
Lemma 4.6.
Let $(Y,D,L)$ be a generic polarized log Calabi–Yau surface, with a polyhedral decomposition $\mathscr{P}$ of a Symington polytope $P$ defined as above. Then, the normal crossing surface $\mathcal{X}_{\mathscr{P},0}$ is a quasi-projective open Kulikov surface.
Proof.
By construction, the dual intersection complex of $\mathcal{X}_{\mathscr{P},0}$ is $(P,\mathscr{P})$. Since $P$ is a disk, condition i) in Definition 3.7 is satisfied. By Lemmas 4.4 and 4.5, the irreducible components of $\mathcal{X}_{0}$ are 0-, 11-, and $2$-surfaces, and so condition ii) in Definition 3.7 holds as well. Finally, since $\mathscr{P}$ is a triangulation into triangles of size one, $\mathcal{X}_{0}$ is combinatorially $d$-semistable as in condition iii) of Definition 3.7 by [32, Proposition 3.6]. Thus, the normal crossing surface $\mathcal{X}_{0}$ is an open Kulikov surface.
To prove quasi-projectivity, recall that, by Definition 2.15, the good polyhedral decomposition $\overline{\mathscr{P}}$ is regular. Consequently, there exists a convex piecewise linear function $\varphi$ on $\overline{P}$ that induces $\overline{\mathscr{P}}$. Let $L_{\varphi}$ denote the corresponding ample line bundle on $\mathcal{X}_{\overline{\mathscr{P}}}$. From the toric description of the deformation in Step I of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, the restrictions of $L_{\varphi}$ to $\bigcup_{v^{\prime}\in pr^{-1}(v)}X^{v}_{\overline{\mathscr{P}}}$ deform into ample line bundles $\widetilde{L}_{\varphi}^{v}$ on the toric surfaces $\widetilde{X}^{v}_{\mathscr{P}}$. Since any toric surface $X$ satisfies $H^{2}(X,\mathcal{O}_{X})=0$, deformation theory for line bundles (see for example the proof of [89, Theorem 2.5.23 (ii)]) ensures that, in Step II of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, the ample line bundles $L^{v}_{\varphi}$ on the toric surfaces $\widetilde{X}^{v}_{\mathscr{P}}$ deform into ample line bundles $L^{v}_{\varphi}$ on the log Calabi–Yau surfaces $X^{v}_{\mathscr{P}}$. Furthermore, for every edge $e\in\mathscr{P}^{[1]}$ adjacent to two vertices $v,v^{\prime}\in\mathscr{P}^{[1]}$, the line bundles $L^{v}_{\varphi}$ and $L^{v^{\prime}}_{\varphi}$ are isomorphic when restricted to $X^{e}_{\mathscr{P}}$. Thus, by the description of $\mathrm{Pic}(\mathcal{X}_{\mathscr{P},0})$ in Lemma 3.10, the ample line bundles $L^{v}_{\varphi}$ glue together to form a global ample line bundle on $\mathcal{X}_{\mathscr{P},0}$. ∎
4.3. Semistable mirrors of polarized log Calabi–Yau surfaces
Let $(Y,D,L)$ be a generic polarized log Calabi–Yau surface and $(P,\mathscr{P})$ a Symington polytope with polyhedral decomposition defining a maximal degeneration of $(Y,D,L)$ as in §2.4. By Theorem 3.31, up to a locally trivial deformation, we can assume that the open Kulikov surface $\mathcal{X}_{\mathscr{P},0}$ is $d$-semistable and generic. Therefore, by Theorem 3.19, there exists a smoothing $\pi_{\mathscr{P}}:\mathcal{X}_{\mathscr{P}}\to\Delta$ of $\mathcal{X}_{\mathscr{P},0}$, which in addition is an open Kulikov degeneration. From now on we refer to this smoothing as the semistable mirror to $(Y,D,L)$.
Although a smoothing of $\mathcal{X}_{\mathscr{P},0}$ is not unique, it is unique up to deformation. Throughout the remainder of this paper, we will always consider smoothings of $\mathcal{X}_{\mathscr{P},0}$ up to deformation. Therefore, to simplify notation, we will call $\mathcal{X}_{\mathscr{P}}$ “the” smoothing of $\mathcal{X}_{\mathscr{P},0}$, and refer to it as the semistable mirror to the polarized log Calabi–Yau surface $(Y,D,L)$, with given Symington polytope $P$ endowed with a good polyhedral decomposition $\mathscr{P}$. In the following section §5, we extend the intrinsic mirror symmetry construction of Gross–Siebert to encompass quasi-projective open Kulikov degenerations such as $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$. Ultimately, we show that starting with a polarized log Calabi–Yau $(Y,D,L)$, constructing its semistable mirror, and then taking the intrinsic mirror to the semistable mirror indeed yields a family of polarized log Calabi–Yau’s deformation equivalent to the initial $(Y,D,L)$, as shown in §8. Thus, the semistable mirror construction can be viewed as the reverse of a generalized intrinsic mirror construction of Gross–Siebert for families of log Calabi–Yau’s.
@@ 4.3. Semistable mirrors of polarized log Calabi–Yau surfaces
Consider the case where $(Y,D)$ is a toric pair, where $Y$ is a toric surface and $D$ is its anti-canonical boundary divisor, and denote by $P$ the momentum polytope corresponding to the ample line bundle $L$ on $Y$. Then, the affinization $\mathcal{X}^{\mathrm{can}}_{\mathscr{P}}\rightarrow\Delta$ of the semistable mirror is the completion along $0\in\mathbb{A}^{1}$ of the three dimensional Gorenstein affine toric variety with fan the cone over $P$ and which admits a natural toric morphism to $\mathbb{A}^{1}$. The semistable mirror $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ is the completion along 0 of the toric crepant resolution of this affine toric singularity defined by the regular triangulation $\mathscr{P}$ of $P$ into triangles of size one.